Cost-Volume-Profit Analysis

Gannon Company sells a single product for $18.75 per unit. Variable costs are $9 per unit and fixed costs are $134,940 at an operating level of 12,000 to 25,000 units.

Round UP answers to the nearest unit, when applicable.

a. What is Gannon Company's break-even point in units? Answer

units

b. How many units must be sold to earn $20,000 before income tax? Answer

units

c. How many units must be sold to earn $30,000 after income tax, assuming a 40% tax rate? Answer

units
Answer #1

Contibution Margin Per Unit = Selling Price - Variable cost Per Unit

CM Per Unit = 18.75 -9.00 = 9.75

A) Break Even Units = Fixed cost / CM Per Unit = 134,940 /9.75 = 13,840 (Answer)

B) Units to get 20,000 Before Tax income = (134940+20000) /9.75 = 15,891.28 = 15,891 Units (Answer)

C) Units To Get 30,000 After Tac income = ( 134,940 +(50,000/0.6) )/9.75 = 18,968 (Answer)